Given A(-5,-4) and B(7,18), find the point two-fifths of the way from A to B.
Found 2 solutions by Alan3354, robertb:
Answer by Alan3354(69443)

Given A(-5,-4) and B(7,18), find the point two-fifths of the way from A to B.
----------
Find the x & y separately.
For x:
Diff = 12
(2/5)*12 = 4.8
-5 + 4.8 = -0.2
=================
For y:
Diff = 22
(2/5)*22 = 8.8
-4 + 8.8 = 4.8
=========================
Point is (-0.2,4.8)
Or (-1/5,24/5)

Answer by robertb(5830)